OpenStudy (anonymous):

Solve the following equation by using the substitution method. Write your answer as an order pair. Y + 3x = -1 2y -5x = 31

OpenStudy (kropot72):

Multiply (y + 3x = -1) by 2. Then subtract the result from (2y - 5x = 31). This will eliminate y and enable you to find the value of x. Can you try this?

OpenStudy (anonymous):

Thanks So Much :)

OpenStudy (muhammad_nauman_umair):

as y=31-3x then -12(31-3x)-5x=31 -372+36x-5x=31 31x=372+31 x=13 and y=-8 (13,-8)

OpenStudy (anonymous):

Thanks =)

OpenStudy (kropot72):

My tested result is (-3,8) for the ordered pair

OpenStudy (kropot72):

Try testing (13,-8) in (y + 3x = -1)